An observer sees a flower pot sail up and then back down past a window 2.45 m high. If the total time the pot is in the sight is 1 s, what is the height above the window upto which the pot rises ?


 

Grade:

1 Answers

AJIT AskiitiansExpert-IITD
13 years ago

Dear  Fawz naim ,

THe time the pot takes to go up and down through the window must be same as the conditions in both cases are same. The magnitude of velocity at the entry and exit points are same and distance travelled is same too . so time taken to go up and down = 0.5 s

now we can take the axis of reference anywhere on the trajectory of the motion , we take it on the bottom of window. Assuming the velociyt at that point to be u , and applying the equaitons of motion ,

2.45 = u*(.5) -(1/2) *9.8*.52

solving we get u = 7.35m/s

so the maximum height from the bottom of the window = u2/2g  = 2.7625

height from the top of window is  2.7625- 2.45 = .306 m

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

Askiitians Expert

Ajit Singh Verma IITD